comelete the following sentences: (20m) l in order for an oscillator to work, the feedback must be - 21 The phase shift of an RC lead-lag network at te is 3) The feedback fraction in a Wien bridge oscilator is 4) A Hartley oscillator has a tapped 5) Crystals are more stable than LC tanks due to their very high 6) The cutoff level of the amplifier can be calculated from the maximum level by using the equation- in its tank cin 7) The cross-over distortion of the power ampli 8) A UIT relaxation oscllator provides t fier is caused by two waveforms, they are 9) The oscillation frequency of Colpitts scilator can be determined using the following expression 10) The abbreviation (Aue) stands for 11) The power dissipation of the power amplfier is the power which is converted to of the power amplifier is the power 12) The tem Power Conversion Efficiency n can be defined as 13) ideal op-amp has input resistance of 14)If 12/1 1 decade, and 1 10 KH, then 12 15) THO (total harmonic distortion) al harmonic distortion) can be calkulated ollowinge ollowing expre 16) The output t resistance of signal source Ru, the input coupling capacitor Ccl, the input of the ampifier Ri form a filter,the filter type is I You have an amplifier with (fu 35 H,f70H24 otul. The bandwidth 8W wil 18) The location of Q point of class B power amplifier is 19) This equation charactenize the operation of class AB power amplfien Explain the meaning linterpret atioe) of this 20) The ideal parameters of op-amp are
Answer #1

1. Be positive and amplitude should be stable.

2. 00

3. 1/3

4. Resonant coil

5. Frequency

6. 20 log (V2/ V1)

7. Signal changes

8. Capacito waveform, sawtoothed waveform

9. fr = 1/ 2\pi\sqrtLCT     Where CT = 1/ C1 + 1/C2

10. Operational amplifier

11. AC voltage signal

12. (Output power in Ac/ Input Power in Dc) X 100

13. Infinite

14. f2 = 10 KHz X 1 Decade

= (10 X 103) Hz X 10

= 105Hz

15. THD = VRMS without fundamental / VRMS fundamental Where VRMS = Root mean speed

16. High pass filter

17. Bw = fh - fl ( highest frequency - lowest frequency)

= ( 5.2 KHz - 35 Hz)

18. at cut- off

19. Very small bias voltage in the two transistor either Bi polar of FET will be on for more than one half cycle but much less than one full cycle of input signal.

20. (i) Infinte voltage gain

(ii) Infinite input impendance

(iii) Zero output impendance

(iv) infinite open loop

(v) closed bandwidth

(vi) infinite slew rate

( vii) infinite CMRR

  (viii) Zero offset

  (ix) No noise